An economy can be described by the production function, ๐‘Œ = ๐น(๐พ, ๐ฟ) = ๐พ ๐›ผ๐ฟ 1โˆ’๐›ผ (a) Show that this production function exhibits constant returns to scale?(b) What is the per-worker production function? (c) Assuming a version of the Solow growth model with population growth but no technological progress, find expressions for the steady-state capital-output ratio, capital stock per worker, output per worker, and consumption per worker, as a function of the saving rate (๐‘ ), the depreciation rate (๐›ฟ), and the population growth rate (๐‘›). (You may assume the condition that capital per worker evolves according to โˆ†๐‘˜ = ๐‘ ๐‘“(๐‘˜) โˆ’ (๐‘› + ๐›ฟ)๐‘˜.) [4 marks] Now consider a specific economy described by the production function, ๐‘Œ = ๐น(๐พ, ๐ฟ) = ๐พ 0.6๐ฟ 0.4 The economy has no technological progress and has a depreciation rate of 5% per year. The economy starts in a steady-state with growth in output (๐‘Œ) of 5% per year. Further, the economy exhibits a capital-output ratio of 2 in this steady-state.


1
Expert's answer
2021-03-24T20:50:44-0400
"solution"


A]constant return to scale


"Y = f (K,L)\\
\frac{Y}{L}=[\frac{1}{l}]F(k,l)=F(\frac{k}{l},\frac{l}{l})=F(\frac{k}{l},1)\\"


"Define\ Y ≡ Y /L \ and\ k ≡ K/L.\\ Then:\\
\ Y = F (k,1) = f (k)\\
Y=F(K,L)"


per-worker

"F (zK, zL) = A(zK)
^α
(zL)
^{1−α}\\
F= z
^α
z
^{1−α}AK^α
L
^{1−α}\\
F= zAK^α
L
^{1−α}\\
F= zF (K,L)"



Steady state

"ss:\Delta K=0\\
Investment\ 
 equals \ depreciation"


"\Delta k=sf(k)-(n-\delta)k"

"Y=F(K,L)"

"0.05y=0.05(0.6\times 0.4\\
0.05y=0.012\\
y=\frac{0.012}{0.05}\\
y=4.166"